# Cosmological Inflation and Dark Sector from 11D Supergravity

### v2

@@ 2.2 General Case
Since we anticipate the presence of Einstein’s Gravity in a four-dimensional spacetime, the 11th dimension is compacted into a circular shape as
@@
with the indices $i,j=1,2,3$. Analogous to the special case presented in Subsection 2.1, the four-dimensional effective action for cosmological D3-branes can be derived through the imposition of appropriate gauge conditions and boundary conditions. Unsurprisingly, the effective action can be written as
$S_{\text{brane}}=\int d^{4}x\sqrt{-g_{4}}\left[\frac{m_{p}^{2}}{2}R_{(4)}-\frac{1}{2}\frac{\lambda m_{p}^{2}}{\Phi^{2}}\nabla_{\mu}\Phi\nabla^{\mu}\Phi-U(\Phi)+\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}\right].$ (62)
By utilizing the technique demonstrated in equations (47) and (48), which imposes specific Dirichlet boundary conditions for the cosmological D3-brane, the potential can be defined as follows:
@@
The corresponding Dirichlet boundary conditions can be derived from the equations
$R_{T^{2}_{q_{i}}}[c_{a_{i}}]=\frac{8\pi(1-q_{i})}{A_{T^{2}_{q_{i}}}},i=1,2,3.$ (64)
@@
To compare our findings with the discussions presented in subsection 2.1, the equations written as $2A_{T_{q_{1}}^{2}}=A_{T_{q_{2}}^{2}}=A_{T_{q_{2}}^{2}}=A_{0}$ remain applicable. As the inflaton $\phi$ rolls down to a true vacuum, we can establish conditions $U(\Phi_{min})=0,U^{\prime}(\Phi_{min})=0$ on the potential. The solution is given by
@@
with $\Phi=-e^{\sqrt{\lambda^{-1}}\phi/m_{p}}$ and $n,q_{3}\in\mathbb{N}$. The pictures of this potential are shown in Figure 1. It can be seen from the listed potential functions that the lowest point of potential energy is located at $U(\Phi_{min})=0$. When $\phi\gg 1$, the potential energy tends to be a flat potential, which has the characteristics of the potential function of the Starobinsky model. An important property of the potential $U(\Phi)$ is that the spectral index $n_{s}$ and tensor-to-scalar ratio $r$ are independent of the factor $(q_{3}-1)$ and parameter $n$. Therefore, the only constraint imposed on the third torus is $q_{3}\geq 2$. For more details, please refer to the Appendix. Moreover, we rediscover the special case in Subsection 2.1 by choosing $\Phi_{min}=1,s_{1}=\frac{1}{2},s_{2}=0,s_{3}=1,q_{1}=0,q_{2}=q_{3}=2$.
@@
3 Cosmological Inflation
3.1 Numerical Results
In this section, we aim to numerically compute the inflation parameters corresponding to the inflation models discussed in Section 2. The inflaton $\phi$ is clearly associated with the size function $\Phi(x^{\mu})$ of the extra dimensions. Therefore, the vibrations of the inflaton $\phi$ can be considered as the conformal vibrations of the extra dimensions.
@@ 3.1 Numerical Results
Figure 2: The $n_{s}-r$ constraints for potential (71). The green contours are from Planck 2018 data Planck:2018jri . The blue contours are from 2021 BICEP/Keck data BICEP:2021xfz . Here, we choose $s_{3}=1,2,3,4,7,10,...,37,40$. The straight lines of varying colors represent distinct values of $s_{3}$. The lowest red straight line represents $s_{3}=1$. As $s_{3}$ increases, the corresponding $n_{s}-r$ line progressively shifts beyond the blue observational region (See the yellow region.). However, as $s_{3}$ continues to grow, these lines tend to cluster around the purple line at the top.
@@
where $\phi_{i}$ represents the initial value of the inflaton and $\phi_{e}$ represents its end value. The numerical results of the general case with potential (71) are shown in Figure 2. The lowest three lines in Figure 2 correspond to the cases of $s_{3}=1,2,3$, respectively. They are consistent with the stronger $n_{s}-r$ constraints by Planck TT,TE,EE+lowE+lensing+BK18+BAO (blue region). With the increasing $s_{3}$, the convergence of $n_{s}-r$ happens inside the green region of weak $n_{s}-r$ constraints. As a result, the inflation models from the compactification methods $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$ can be in good accordance with the latest observations. The optimal value for $s_{3}$ is $s_{3}=1$, as its corresponding $n_{s}-r$ line approaches the center of the observational constraints, which is the same as the result of the special case discussed in Subsection 2.1.
For the special case (53), the inflation exits as $\epsilon(\phi_{e})=1,\phi_{e}=1.182m_{p}$. For e-folding $N=60$, the solutions are $(\phi_{i},\epsilon,\eta,n_{s},r)=(8.986m_{p},0.0010,-0.012,0.9695,0.016)$. For e-folding $N=50$, $(\phi_{i},\epsilon,\eta,n_{s},r)=(8.503m_{p},0.0013,-0.014,0.9637,0.022)$. These results match the Planck 2018 data Planck:2018jri . Taking into account the observational constraint imposed on the power spectrum Planck:2018vyg
@@
It shows that the scale of the extra dimensions is considerably small at order $O(10^{5}l_{p})$. In our compactification methods, the eleventh dimension is compactified into a circle, indicating that the theory resulting from compactification is Type IIA string theory in 10 dimensions. Subsequently, the string length $l_{s}$ relates to the radius of the eleventh dimension $\mathcal{R}_{10}$ as follows Ydri:2018sjn :
@@
The standard tori extra dimensions have a size scale of $\mathcal{R}=29387l_{p}=3.39\xi^{\frac{1}{4}}l_{p(11)}$. Assuming the loop-order quantum corrections of eleven-dimensional supergravity Hyakutake:2006aq can be neglected due to their small coupling constant $\alpha^{\prime}\sim O(l^{2}_{s})$, the string length and the size scale of the tori must satisfy the condition $\frac{l_{s}}{\mathcal{R}}\ll 1$. This limitation imposes a constraint on the parameter $\xi\ll 0.025$. Therefore, we anticipate that the current string length meets the criterion $l_{s}=8676.7\xi^{\frac{3}{4}}l_{p}\ll 551.9l_{p}$.
The dynamics of the inflaton and the scale function are shown in figure 3. It demonstrates that the three tori contract conformally, while the eleventh dimension expands. The scale function oscillates around the point where $\Phi(x^{\mu})$ is equal to 1 after the process of inflation.
3.2 Exit Inflation
There exist additional matter terms within the four-dimensional effective action of 11D Supergravity (41). Because the 11D metric $G_{AB}$ couples to gravitino $\psi_{A}$ and 3-form $A_{MNP}$, the inflaton $\phi$, associated to the size function $\Phi$ of extra dimensions, will couple to some components of $\psi_{A}$ and $A_{MNP}$ after compactification. When the inflaton $\phi$ begins to decay into the gravitino and the 3-form, the inflation exits. In this period, we keep the gauge condition $\delta g_{\mu\nu}=0$ and release $\delta g_{a_{1}b_{1}},\delta h_{a_{2}b_{2}},\delta\gamma_{a_{3}b_{3}}\neq 0$. This decay process gives rise to a non-zero $\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}$ term in the four-dimensional effective action, as seen in equation (44). For example, matter in the universe can be generated by introducing the interactions between the inflaton and the Higgs field $H$, expressed as $\xi\phi^{2}H^{2}$, into the Lagrangian $\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}$.
4 Dark Sector from Extra Dimensions
In this section, we concentrate on the quantum effects from the inflaton in the four-dimensional spacetime specifically for the case presented in Subsection 2.1. As shown in Equation (80), the extra dimensions exist at the sub-Planck scale $O(10^{5}l_{p})$, and their conformal vibrations can be interpreted as the vibrations of the inflaton $\phi$. Referring to Figure 3, it becomes evident that the inflaton field arrives at a significantly small value, namely $\phi\ll m_{p}$, in today’s universe. Hence, the potential given in (54) can be expanded up to the second order as
@@ 4 Dark Sector from Extra Dimensions
Following the second quantization method, the inflaton field emerges as a massive scalar quantum field. The tree-level propagator for graviton with momentum $k^{\mu}=(k_{0},\mathbf{k})$ is given by Mohanty:2022abo
@@
where the Minkowski metric $\eta_{\mu\nu}=\text{diag}(-1,1,1,1)$ in the local reference frames. Consider the action (53) along with the matter term $\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}$, the energy stress tensor can be represented as
@@
where the energy stress tensor of the inflaton $\mathcal{T}^{\mu\nu}_{\phi}$, as well as the density $\rho$ and the pressure $p$, are defined by
@@
$\rho$ $=\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}\partial_{i}\phi\partial^{i}\phi+V(\phi),$ (95)
$p$ $=\frac{1}{2}\dot{\phi}^{2}-\frac{1}{2}\partial_{i}\phi\partial^{i}\phi-V(\phi).$ (96)
The inflaton fields can now be regarded as a form of dark energy for $\rho+3p<0$, or cold dark matter for $p/\rho\approx 0$, pervading the entire spacetime. In classical mechanics, when considering two massive point particles with masses $M_{1}$ and $M_{2}$, Newton’s gravity law has modification due to the interactions from the inflaton. The four-point tree-level scattering amplitude for the matter-graviton-matter process can be expressed as the effective action of the second order Mohanty:2022abo
@@
The four-vector $u^{\mu}=(1-\mathbf{v}^{2})^{-\frac{1}{2}}(1,\mathbf{v})=I^{\mu}+O(\mathbf{v})$, where $I^{\mu}=(1,0,0,0)$. This four-point tree-level scattering amplitude contains both the Newton potential part between $M_{1}$ and $M_{2}$, denoted as Mohanty:2022abo
@@
and the self-gravity interaction term of inflation, denoted as
@@ 4.1 Dark Matter
Now, setting $\mathbf{q}_{1}(t)=0$ and locating the center of the dark matter distribution at the origin $\mathbf{q}_{1}(t)$, we follow the method in Ref. Fitzpatrick_2012 to calculate Eq. (105). Hence Newton’s law of gravity for $M_{2}$ is given as follows
@@
The characteristic length of the Yukawa term, given by $\lambda=\frac{\hbar}{2am_{2}c}$, constrains the value of $m_{2}$ based on experimental results on Yukawa gravitational potential Tan:2020vpf ; Lee:2020zjt and observation results on galaxies. Considering $m_{2}$ correction, when $m_{2}R\gg 1$, the velocity decays in the manner of $v_{2}^{2}\propto R^{-1}$ which reflects the original Newton’s gravity law still holds at an extremely large distance but additional mass from dark matter to $M_{1}$ should be included as the central gravitational source. When $m_{2}R\ll 1$, Eq. (113) in the case of $C_{1}=0$ is recovered, and the velocity of $M_{2}$ is directly expressed by equation
@@
Equation (122) recovers the formulation of the MOND theory Milgrom:1983zz ; Milgrom:1983ca ; Bekenstein:1984tv ; Milgrom:1986ib . Then, we find the Tully-Fisher relation from (119)(123) as $R\rightarrow\infty$
@@
These are two of the ways in which extra dimensions affect Newton’s gravity law. In this scenario, the extra dimensions are viewed as a spherically symmetric distribution of dark matter. Astronomical observations, as referenced in Das:2023zvp , reveal that the characteristic acceleration is of the order of $a_{0}\sim 10^{-8}[\text{cm/}\text{s}^{2}]$ and the radii of galaxies are approximately $R_{\text{max}}\sim 10^{2}[\text{kpc}]\approx 3.8\times 10^{55}l_{p}$. Consequently, the characteristic mass becomes $m_{2c}=1/R_{\text{max}}\approx 2.6\times 10^{-56}m_{p}$. Therefore, we observe the relations $m_{1}\gg m_{2c}\gg m_{2}$. The modified portion of Newton’s gravity law is associated with both the mass of the inflaton $m_{\phi}$ and the strength of the inflaton field
@@
It is worth noting that certain galaxies, as observed in astronomy Comeron:2023dop ; PinaMancera:2021wpc ; vanDokkum:2018vup ; vanDokkum:2022zdd ; Danieli:2019zyi , lack dark matter. An explanation for this phenomenon could be that their characteristic acceleration $a_{0}$ and mass $M_{1}$ are small enough to result in $C^{2}\sim 0$. In such galaxies, the inflaton distribution appears to be relatively uniform and sparse so that the scale function of extra dimensions $\Phi(x^{\mu})\approx 1$. The relativistic contribution at the leading order to Newton’s gravity law discussed in this subsection differs from the one presented in Das:2023zvp . The results outlined in (105) indicate an effective dark matter density profile
@@ 4.2 Dark Energy
This solution suggests that the extra dimensions can be interpreted as a spherically symmetric distribution of dark energy. This result aligns with the experimental constraints on the Yukawa gravitational potential Tan:2020vpf ; Lee:2020zjt , where the characteristic length $\lambda=\frac{\hbar}{2m_{\phi}c}\approx 4.04\times 10^{-30}[\text{m}]$, but it is far beyond the current experimental accuracy to detect the gravitational effect (132) at centimeter scale. It’s worth noting that in this scenario, $w=\frac{p_{\text{DE}}}{\rho_{\text{DE}}}\approx-1$ indicating that the inflaton might be just one component of dark energy. However, the effective action (44) represents the tree-level effective action for eleven-dimensional supergravity. Higher-order derivatives of the inflaton, such as $\nabla^{\mu}\nabla^{\nu}{\phi}\nabla_{\mu}\nabla_{\nu}{\phi}$, can be found in the 1-loop order action of eleven-dimensional supergravity. We leave further discussions on the role of the inflaton as dark energy for future study.
5 Conclusions and Discussion

### v3

@@ II.2 General Case
Since we anticipate the presence of Einstein’s Gravity in a four-dimensional extended spacetime, the 11th dimension is compacted into a circular shape as
@@
with the indices $i,j=1,2,3$. Analogous to the special case presented in Subsection II.1, the four-dimensional effective action for 4D extended spacetime can be derived through the imposition of appropriate gauge conditions and boundary conditions. Unsurprisingly, the effective action of the 11D bulk can be written as
$S_{\text{eff}}=\int d^{4}x\sqrt{-g_{4}}\left[\frac{m_{p}^{2}}{2}R_{(4)}-\frac{1}{2}\frac{\lambda m_{p}^{2}}{\Phi^{2}}\nabla_{\mu}\Phi\nabla^{\mu}\Phi-U(\Phi)+\mathcal{L}_{\text{matter}}\right].$ (37)
By utilizing the technique demonstrated in equations (23) and (24), the potential can be defined as follows:
@@
To compare our findings with the discussions presented in subsection II.1, the equations written as $2A_{T_{q_{1}}^{2}}=A_{T_{q_{2}}^{2}}=A_{T_{q_{2}}^{2}}=A_{0}$ remain applicable. As the inflaton $\phi$ rolls down to a true vacuum, we can establish conditions $U(\Phi_{min})=0,U^{\prime}(\Phi_{min})=0$ on the potential. The solution is given by
@@
with $\Phi=-e^{\sqrt{\lambda^{-1}}\phi/m_{p}}$ and $n,q_{3}\in\mathbb{N}$. The pictures of this potential are shown in Figure 1. It can be seen from the listed potential functions that the lowest point of potential energy is located at $U(\Phi_{min})=0$. When $\phi\gg 1$, the potential energy tends to be a flat potential, which has the characteristics of the potential function of the Starobinsky model. An important property of the potential $U(\Phi)$ is that the spectral index $n_{s}$ and tensor-to-scalar ratio $r$ are independent of the factor $(q_{3}-1)$ and parameter $n$. Therefore, the only constraint imposed on the third torus is $q_{3}\geq 2$. For more details, please refer to the B. Moreover, we rediscover the special case in Subsection II.1 by choosing $\Phi_{min}=1,s_{1}=\frac{1}{2},s_{2}=0,s_{3}=1,q_{1}=0,q_{2}=q_{3}=2$.
@@
III Cosmological Inflation
In this section, we aim to numerically compute the inflation parameters corresponding to the inflation models discussed in Section II. The inflaton $\phi$ is clearly associated with the size function $\Phi(x^{\mu})$ of the extra dimensions. Therefore, the vibrations of the inflaton $\phi$ can be considered as the conformal vibrations of the extra dimensions.
@@ III Cosmological Inflation
Figure 2: The $n_{s}-r$ constraints for potential (45). The green contours are from Planck 2018 data [2]. The blue contours are from 2021 BICEP/Keck data [1]. Here, we choose $s_{3}=1,2,3,4,7,10,...,37,40$. The straight lines of varying colors represent distinct values of $s_{3}$. The lowest red straight line represents $s_{3}=1$. As $s_{3}$ increases, the corresponding $n_{s}-r$ line progressively shifts beyond the blue observational region (See the yellow region.). However, as $s_{3}$ continues to grow, these lines tend to cluster around the purple line at the top.
@@
where $\phi_{i}$ represents the initial value of the inflaton and $\phi_{e}$ represents its end value. The numerical results of the general case with potential (45) are shown in Figure 2. The lowest three lines in Figure 2 correspond to the cases of $s_{3}=1,2,3$, respectively. They are consistent with the stronger $n_{s}-r$ constraints by Planck TT,TE,EE+lowE+lensing+BK18+BAO (blue region). With the increasing $s_{3}$, the convergence of $n_{s}-r$ happens inside the green region of weak $n_{s}-r$ constraints. As a result, the inflation models from the compactification methods $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$ can be in good accordance with the latest observations. The optimal value for $s_{3}$ is $s_{3}=1$, as its corresponding $n_{s}-r$ line approaches the center of the observational constraints, which is the same as the result of the special case discussed in Subsection II.1.
For the special case (29), the inflation exits as $\epsilon(\phi_{e})=1,\phi_{e}=1.182m_{p}$. For e-folding $N=60$, the solutions are $(\phi_{i},\epsilon,\eta,n_{s},r)=(8.986m_{p},0.0010,-0.012,0.9695,0.016)$. For e-folding $N=50$, $(\phi_{i},\epsilon,\eta,n_{s},r)=(8.503m_{p},0.0013,-0.014,0.9637,0.022)$. These results match the Planck 2018 data [2]. Taking into account the observational constraint imposed on the power spectrum [59]
@@
It shows that the scale of the extra dimensions is considerably small at order $O(10^{5}l_{p})$. In our compactification methods, the eleventh dimension is compactified into a circle, indicating that the theory resulting from compactification is Type IIA string theory in 10 dimensions. Subsequently, the string length $l_{s}$ relates to the radius of the eleventh dimension $\mathcal{R}_{10}$ as follows [60]:
@@
The standard tori extra dimensions have a size scale of $\mathcal{R}=29387l_{p}=3.39\xi^{\frac{1}{4}}l_{p(11)}$. Assuming the loop-order quantum corrections of eleven-dimensional supergravity [61] can be neglected due to their small coupling constant $\alpha^{\prime}\sim O(l^{2}_{s})$, the string length and the size scale of the tori must satisfy the condition $\frac{l_{s}}{\mathcal{R}}\ll 1$. This limitation imposes a constraint on the parameter $\xi\ll 0.025$. Therefore, we anticipate that the current string length meets the criterion $l_{s}=8676.7\xi^{\frac{3}{4}}l_{p}\ll 551.9l_{p}$.
The dynamics of the inflaton and the scale function are shown in figure 3. It demonstrates that the three tori contract conformally, while the eleventh dimension expands. The scale function oscillates around the point where $\Phi(x^{\mu})$ is equal to 1 after the process of inflation, see C for a brief discussion on exit inflation.
IV Dark Sector from Extra Dimensions
In this section, we concentrate on the quantum effects from the inflaton in the four-dimensional extended spacetime specifically for the case presented in Subsection II.1. As shown in Equation (54), the extra dimensions exist at the sub-Planck scale $O(10^{5}l_{p})$, and their conformal vibrations can be interpreted as the vibrations of the inflaton $\phi$. Referring to Figure 3, it becomes evident that the inflaton field arrives at a significantly small value, namely $\phi\ll m_{p}$, in today’s universe. Hence, the potential given in (30) can be expanded up to the second order as
@@ IV Dark Sector from Extra Dimensions
Following the second quantization method, the inflaton field emerges as a massive scalar quantum field. The tree-level propagator for graviton with momentum $k^{\mu}=(k_{0},\mathbf{k})$ is given by [62]
@@
where the Minkowski metric $\eta_{\mu\nu}=\text{diag}(-1,1,1,1)$ in the local reference frames. Consider the action (29) along with the matter term $\mathcal{L}_{\text{matter}}$, the energy stress tensor can be represented as
@@
where the energy stress tensor of the inflaton $\mathcal{T}^{\mu\nu}_{\phi}$, as well as the density $\rho$, are defined by
@@
$\rho$ $=\frac{1}{2}\dot{\phi}^{2}+\frac{1}{2}\partial_{i}\phi\partial^{i}\phi+V(\phi).$ (69)
In classical mechanics, when considering two massive point particles with masses $M_{1}$ and $M_{2}$, Newton’s gravity law has modification due to the interactions from the inflaton. The four-point tree-level scattering amplitude for the matter-graviton-matter process can be expressed as the effective action of the second order [62]
@@
The four-vector $u^{\mu}=(1-\mathbf{v}^{2})^{-\frac{1}{2}}(1,\mathbf{v})=I^{\mu}+O(\mathbf{v})$, where $I^{\mu}=(1,0,0,0)$. This four-point tree-level scattering amplitude contains both the Newton potential part between $M_{1}$ and $M_{2}$, denoted as [62]
@@
and the self-gravity interaction term of inflaton, denoted as
@@ IV.1 Dark Matter
Now, setting $\mathbf{q}_{1}(t)=0$ and locating the center of the dark matter distribution at the origin $\mathbf{q}_{1}(t)$, we follow the method in Ref. [63] to calculate Eq. (78). Hence Newton’s law of gravity for $M_{2}$ is given as follows
@@
The characteristic length of the Yukawa term, given by $\lambda=\frac{\hbar}{2am_{2}c}$, constrains the value of $m_{2}$ based on experimental results on Yukawa gravitational potential [64, 65] and observation results on galaxies. Considering $m_{2}$ correction, when $m_{2}R\gg 1$, the velocity decays in the manner of $v_{2}^{2}\propto R^{-1}$ which reflects the original Newton’s gravity law still holds at an extremely large distance but additional mass from dark matter to $M_{1}$ should be included as the central gravitational source. When $m_{2}R\ll 1$, Eq. (86) in the case of $C_{1}=0$ is recovered, and the velocity of $M_{2}$ is directly expressed by equation
@@
Equation (95) recovers the formulation of the MOND theory [43, 44, 45, 46]. Then, we find the Tully-Fisher relation from (92)(96) as $R\rightarrow\infty$
@@
These are two of the ways in which extra dimensions affect Newton’s gravity law. In this scenario, the extra dimensions are viewed as a spherically symmetric distribution of dark matter. Astronomical observations, as referenced in [66], reveal that the characteristic acceleration is of the order of $a_{0}\sim 10^{-8}[\text{cm/}\text{s}^{2}]$ and the radii of galaxies are approximately $R_{\text{max}}\sim 10^{2}[\text{kpc}]\approx 3.8\times 10^{55}l_{p}$. Consequently, the characteristic mass becomes $m_{2c}=1/R_{\text{max}}\approx 2.6\times 10^{-56}m_{p}$. Therefore, we observe the relations $m_{1}\gg m_{2c}\gg m_{2}$. The modified portion of Newton’s gravity law is associated with both the mass of the inflaton $m_{\phi}$ and the strength of the inflaton field
@@
It is worth noting that certain galaxies, as observed in astronomy [67, 68, 69, 70, 71], lack dark matter. An explanation for this phenomenon could be that their characteristic acceleration $a_{0}$ and mass $M_{1}$ are small enough to result in $C^{2}\sim 0$. In such galaxies, the inflaton distribution appears to be relatively uniform and sparse so that the scale function of extra dimensions $\Phi(x^{\mu})\approx 1$. The relativistic contribution at the leading order to Newton’s gravity law discussed in this subsection differs from the one presented in [66]. The results outlined in (78) indicate an effective dark matter density profile
@@ IV.2 Dark Energy
This solution suggests that the extra dimensions can be interpreted as a spherically symmetric distribution of dark energy. This result aligns with the experimental constraints on the Yukawa gravitational potential [64, 65], where the characteristic length $\lambda=\frac{\hbar}{2m_{\phi}c}\approx 4.04\times 10^{-30}[\text{m}]$, but it is far beyond the current experimental accuracy to detect the gravitational effect (105) at centimeter scale. It’s worth noting that in this scenario, $w=\frac{p_{\text{DE}}}{\rho_{\text{DE}}}\approx-1$ indicating that the inflaton might be just one component of dark energy. However, the effective action (29) represents the tree-level effective action for eleven-dimensional supergravity. Higher-order derivatives of the inflaton, such as $\nabla^{\mu}\nabla^{\nu}{\phi}\nabla_{\mu}\nabla_{\nu}{\phi}$, can be found in the 1-loop order action of eleven-dimensional supergravity. We leave further discussions on the role of the inflaton as dark energy for future study.
V Conclusions and Discussion